OpenStudy (anonymous):

If F(u) is a homogeneous function of degree n in x and y,then show dat x*(del u/delx)+y*(del u/del y)=n*[F(u)/F'(u)]

OpenStudy (anonymous):

This?\[ x\frac{\partial u}{\partial x}+y\frac{\partial u}{\partial y}=n\frac{F(u)}{F'(u)} \]

OpenStudy (anonymous):

yes this one we have to prove

OpenStudy (anonymous):

Hmmm, may be something like chain rule would help? \[ F'(u)\frac{\partial u}{\partial x}=F_x(u) \]

OpenStudy (anonymous):

I'm not completely sure. Are there any interesting theorems that were introduce recently in the class?

OpenStudy (anonymous):

yes euler's theorem

OpenStudy (anonymous):

dis s a question of partial derivative

OpenStudy (anonymous):

Can you state Euler's theorem?

OpenStudy (anonymous):

if f(x,y,z) be a homogeneous function of degree n,then

OpenStudy (anonymous):